Part II Three vectors A, B, and C are shown below. Question 9. Which of the following equation best describes the relationship between vectors A, B, and C? A BC No equation can be written to relate them Other:

Three vectors A, B, and C are shown.

I am not sure if it C=A+B OR B=A+C? Because both go from tip to tail. Any help would be greatly appreciated!
Answer #1

It's neither C=A+B nor B=A+C because for that C and B would have to be in opposite direction respectively.

Take any of these two case. Let's take the first.

In opposite direction means -C.

So this equation is true:

B+A=-C

=> B+A+C =0

OR A + B + C = 0 (answer)